WARNING: This web is not used anymore. Please use PDBService.MinuteS17July08 instead!
 

Minutes 17 July 08

  • Phone:
    • Gridka: Camile
    • Triumf: Andrew
    • RAL: Carmine
    • BNL: Carlos
    • PIC: Luis

  • CERN:
    • 3D: Dirk, Maria, Luca, Eva
    • ATLAS: Gancho, Florbela

  • Today cannot join - reports sent by email:
    • Barbara from CNAF
    • Alexander from SARA upgrading to 10.2.0.4 at this moment

  • Please review your email lists for the summer period!!

  • Sites status:
    • CERN: IN2P3 databases have been migrated to the new hardware and Streams replication has been reconfigured.
    • GridKA: Nothing to report. Upgrade to 10.2.0.4 foreseen on Sunday 19th and 24th.
    • Triumf: 10.2.0.4 upgrade done
    • RAL: 10.2.0.4 upgrade is not done. Problem with clusterware. Dbs are online but not upgraded. Report will be sent. Try again in few weeks. Adding 3rd node for ATLAS RAC planned to the end of the month.
    • BNL: Nothing to report. Agents state in OEM? State changes to unknown without any reason. SR opened by Carlos.
    • PIC: Upgrade to 10.2.0.4 done!

  • Replication Conditions data to Tier2:
    • Tier1 to Tier2 implications: current setup to be modified, Tier1 sites need to know split and merge
    • Add Tier2 to replication from CERN: is only one site? More Tier2 sites requesting same thing?

  • New licenses: Dirk will send an email asking for team code to pay the licenses money back to CERN. Persons who need to receive the bill can contact Dirk and Maria.

  • Experiments report:
    • ATLAS:
      • TAGs tests Triumf, BNL and PIC.
      • Florbela is interested on I/O measurements with Orion at Triumf, BNL and CERN.
      • Small account (metadata) to be replicated to some sites including DESY. Method to be decided (Streams or other).
      • Carlos is wondering about running Orion in the production system. Luca suggested to use Orion in read only mode. Also, in next intervention this can be done.
      • PVSS replication running for 2 weeks. New several accounts (10 more) will be added to the replication mid August.
      • Total rate about 250 LCRs. Gancho will agree with Eva on the way and time scale.

  • Next meeting – 28th August

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-08-26 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback